# Flow-GRPO: Training Flow Matching Models via Online RL

### v2

@@ Abstract
We propose Flow-GRPO, the first method integrating online reinforcement learning (RL) into flow matching models. Our approach uses two key strategies: (1) an ODE-to-SDE conversion that transforms a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) that matches the original model’s marginal distribution at all timesteps, enabling statistical sampling for RL exploration; and (2) a Denoising Reduction strategy that reduces training denoising steps while retaining the original inference timestep number, significantly improving sampling efficiency without performance degradation. Empirically, Flow-GRPO is effective across multiple text-to-image tasks. For complex compositions, RL-tuned SD3.5 generates nearly perfect object counts, spatial relations, and fine-grained attributes, boosting GenEval accuracy from $63\%$ to $95\%$. In visual text rendering, its accuracy improves from $59\%$ to $92\%$, significantly enhancing text generation. Flow-GRPO also achieves substantial gains in human preference alignment. Notably, very little reward hacking occurred, meaning rewards did not increase at the cost of appreciable image quality or diversity degradation.
@@ 1 Introduction
Flow matching [2, 3] models have become dominant in image generation [4, 5] due to their solid theoretical foundations and strong performance in producing high-quality images. However, they often struggle with composing complex scenes involving multiple objects, attributes, and relationships [6, 7], as well as text rendering [8]. At the same time, online reinforcement learning (RL) [9] has proven highly effective in enhancing the reasoning capabilities of large language models (LLMs) [10, 11]. While previous research has mainly focused on applying RL to early diffusion-based generative models [12] and offline RL techniques like direct preference optimization [13] for flow-based generative models [14, 15], the potential of online RL in advancing flow matching generative models remains largely unexplored. In this study, we explore how online RL can be leveraged to effectively improve flow matching models.
@@
Training flow models with RL presents several critical challenges: (1) Flow models rely on a deterministic generative process based on ordinary differential equations (ODEs) [3], meaning they cannot sample stochastically during inference. In contrast, RL relies on stochastic sampling to explore environments, where the agent learns by trying different actions and receiving feedback. This need for stochasticity in RL conflicts with the deterministic nature of flow matching models. (2) Online RL requires efficient sampling to gather training data, but flow models typically involve many iterative steps to generate each sample, significantly reducing sampling efficiency. This issue becomes even more critical with large, advanced models [5, 4], which are computationally intensive. Therefore, to make RL effective in tasks such as image or video generation, improving sampling efficiency within RL training is crucial.
@@
To address these challenges, we propose Flow-GRPO, which integrates GRPO [16] into flow matching models for text-to-image (T2I) generation, using two key strategies. First, we adopt the ODE-to-SDE strategy to overcome the deterministic nature of the original flow model. By converting the ODE-based flow into an equivalent Stochastic Differential Equation (SDE) framework, we introduce randomness while preserving the original marginal distributions. Second, to improve sampling efficiency in online RL, we apply the Denoising Reduction strategy—reducing denoising steps during training while keeping the full schedule during inference. Our experiments show that using fewer steps maintains performance while significantly reducing data generation costs.
@@
We evaluate Flow-GRPO on T2I tasks with various reward types. (1) Verifiable rewards, using the GenEval [17] benchmark and visual text rendering task. GenEval includes compositional image generation tasks (e.g., generating specific object counts, colors, and spatial relationships), which can be automatically assessed with traditional detection methods. Flow-GRPO improves the accuracy of Stable Diffusion 3.5 Medium (SD3.5-M) [4] from 63% to 95% on GenEval, outperforming the state-of-the-art GPT-4o [18] model. For visual text rendering, SD3-M’s accuracy increases from 59% to 92%, greatly enhancing its text generation ability. (2) Model-based rewards, such as the human preference Pickscore [19] reward, show that our framework is task independent, demonstrating its generalizability and robustness. Importantly, all improvements are achieved with very little reward hacking, as demonstrated in Figure 1.
@@
  • We are the first to introduce GRPO to flow matching models by converting deterministic ODE sampling into SDE sampling, showing that online RL is highly effective for T2I tasks. Specifically, our Flow-GRPO improves SD3.5-Medium accuracy from 63% to 95% without compromising image quality.
  • We find that online RL for flow matching models can use fewer denoising steps for training sample generation, significantly accelerating training. Full denoising steps are still used during testing to maintain performance.
  • We show that the Kullback-Leibler (KL) constraint is effective in preventing reward hacking, where reward increases but image quality or diversity decreases. We also emphasize that KL regularization is not empirically equivalent to early stopping. By adding an appropriate KL, we can achieve the same high reward as the KL-free version while maintaining image quality, though with a longer training time.
@@ 2 Related Work
Online RL has proven effective in boosting the reasoning abilities of large language models (LLMs), notably DeepSeek-R1 [10] and OpenAI-o1 [11]. By iteratively alternating between exploration and exploitation, these systems generate increasingly detailed responses and steadily enhance overall performance. Most successful pipelines employ policy gradient algorithms such as Proximal Policy Optimization (PPO) [20] and the value-free Group Relative Policy Optimization (GRPO) [16]. Because GRPO dispenses with a separate value network, it is more memory efficient; hence we adopt GRPO in this work. Nonetheless, PPO could be incorporated into a flow matching model through an analogous procedure.
@@
Diffusion-based [21, 22, 23] generative models create a forward Markov chain that incrementally adds Gaussian noise to data and train a neural network to approximate the reverse-time denoising vector field; integrating this learned field by discrete DDPM steps or probability flow SDE solvers can yields high fidelity samples. Flow matching [2, 3] instead treats generation as learning a continuous-time normalizing flow whose vector field is trained by directly matching the velocity, enables deterministic sampling in a few ODE steps. Thanks to its efficiency, flow matching attains competitive FID with orders-of-magnitude fewer denoising steps than diffusion samplers and has already become the dominant approach in current image generation [4, 5] and video generation [24, 25, 26, 27] models.  [28, 29] unifies flow-based and diffusion-based methods using the SDE and ODE perspective. Our work builds on their theoretical foundations and introduce GRPO to flow-based models.
@@
Recent work on aligning pretrained T2I models with human preferences has pursued several complementary directions. The main approaches are: (1) direct fine-tuning with scalar reward signals [30, 31, 32, 33]; (2) Reward-Weighted Regression (RWR) [34, 35, 36]; (3) Direct Preference Optimization (DPO) and related variants,[37, 38, 39, 14, 40, 41, 42, 43, 44, 36]; (4) Proximal Policy Optimization (PPO)-style policy gradients [45, 46, 47, 48, 49]; and (5) training-free alignment techniques [50, 51, 52]. These methods have successfully aligned T2I models with human preferences, improving aesthetics and semantic consistency. Building on this progress, we introduce GRPO for flow-matching models—the backbone of today’s state-of-the-art T2I systems. Concurrent work [53] apply GRPO to text-to-speech flow models, but instead of converting the ODE to an SDE to inject stochasticity, they reformulate velocity prediction as estimating a Gaussian distribution (predicting both the mean and variance of velocity), which necessitates retraining the pretrained model. [54] also recognize that stochasticity can be introduced from an SDE perspective, yet their work concentrates on inference-time scaling.
@@ 3 Preliminaries
In this section, we introduce the mathematical formulation of flow matching (Section 3.1) and describe how the denoising process can be mapped as a multi-step MDP (Section 3.2).
3.1 Flow Matching
Let ${\bm{x}}_{0}\sim X_{0}$ denote a data sample drawn from the true data distribution, and ${\bm{x}}_{1}\sim X_{1}$ denote a noise sample. Recent advanced image-generation models (e.g., [4, 5]) and video-generation models (e.g., [24, 26, 25, 27]) adopt the Rectified Flow [3] framework, which defines the “noised” data ${\bm{x}}_{t}$ as
@@ 3.1 Flow Matching
3.2 Denoising as an MDP
According to [12], the iterative denoising process in diffusion models can be framed as an Markov decision process (MDP) $(\mathcal{S},\mathcal{A},\rho_{0},P,R)$. The state at time step p $t$ is ${\bm{s}}_{t}\!\triangleq\!({\bm{c}},t,{\bm{x}}_{t})$, the action is the denoised sample ${\bm{a}}_{t}\!\triangleq\!{\bm{x}}_{t-1}$ produced by the network, and the policy is the conditional model $\pi({\bm{a}}_{t}\mid{\bm{s}}_{t})\!\triangleq\!p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$. The transition kernel is deterministic $P({\bm{s}}_{t+1}\mid{\bm{s}}_{t},{\bm{a}}_{t})\!\triangleq\!(\delta_{\bm{c}},\delta_{t-1},\delta_{{\bm{x}}_{t-1}})$ and the initial-state distribution is $\rho_{0}({\bm{s}}_{0})\!\triangleq\!(p({\bm{c}}),\delta_{T},\mathcal{N}(\mathbf{0},\mathbf{I}))$, in which $\delta_{y}$ is the Dirac delta distribution with nonzero density only at $y$. The reward is non-zero only at the final step, $R({\bm{s}}_{t},{\bm{a}}_{t})\!\triangleq\!r({\bm{x}}_{0},{\bm{c}})$ if $t\!=\!0$ and 00 otherwise.
@@ 4 Flow-GRPO
In this section, we present the Flow-GRPO method, a GRPO-inspired algorithm that improves flow models through online RL. We begin by revisiting the core concept of GRPO [16] and recasting it in the context of flow matching. Next, we discuss how to convert the deterministic probability-flow ODE sampler into a reverse-time SDE with equivalent marginal distribution, thereby introducing stochasticity into the generation process—an essential ingredient for applying GRPO or other RL techniques. Finally, we introduce Denoise Reduction (Section 4.3), a practical data sampling method that dramatically accelerates training without degrading performance.
4.1 GRPO on Flow Matching
Reinforcement learning aims to learn a policy that maximizes expected cumulative reward under a given environment. Typically the goal is to optimize a policy $\pi_{\theta}$ with a regularized objective of the form:
@@ 4.1 GRPO on Flow Matching
Among various RL algorithms, GRPO [16] provides a lightweight alternative to policy based methods such as PPO [20]. Unlike PPO [20], which requires training a separate value function—normally a separate model of similar size to the policy network, GRPO introduces a group relative formulation to estimate the advantage. In our setting, we recast the GRPO framework into the context of flow matching. Recall that the denoising process can be formulated as an MDP, as shown in Section 3.2. Given a prompt ${\bm{c}}$, the flow model $p_{\theta}$ samples a group of $G$ individual images $\{{\bm{x}}^{i}_{0}\}_{i=1}^{G}$ and the corresponding reverse-time trajectories $\{({\bm{x}}^{i}_{T},{\bm{x}}^{i}_{T-1},\cdots,{\bm{x}}^{i}_{0})\}_{i=1}^{G}$. Then, the advantage of the $i$-th image is calculated by normalizing the group-level rewards as follows:
@@
GRPO employs a clipped objective of Eq. 3, together with a KL penalty term:
$\mathcal{J}_{\text{Flow-GRPO}}(\theta)$ $=\mathbb{E}_{{\bm{c}}\sim\mathcal{C},\{{\bm{x}}^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid{\bm{c}})}$ (5)
$\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg{(}\min\Big{(}r^{i}_{t}(\theta)\hat{A}^{i}_{t},\ \text{clip}\Big{(}r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\Big{)}\hat{A}^{i}_{t}\Big{)}-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Bigg{)}\Bigg{]},$
@@
$r^{i}_{t}(\theta)=\frac{p_{\theta}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}{p_{\theta_{\text{old}}}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}.$ (6)
4.2 ODE-to-SDE
In Eq. 4 and Eq. 6, GRPO relies on a stochastic sampling procedure in order to generate a diverse batch of trajectories for advantage estimation and policy exploration. In the case of diffusion models, the forward process consists of a sequence of random Gaussian noising steps, and its reverse-time sampler is implemented via a Markov chain of Gaussians with progressively decreasing variance—equivalently, an approximate solver for the score-based SDE. Thus, diffusion sampling is naturally stochastic.
While in the case of probability flow model, as previously discussed in Eq. 1, its forward process can be formulated as a deterministic ODE:
@@ 4.2 ODE-to-SDE
where ${\bm{v}}_{t}$ is modeled via the flow matching objective in Eq. 2. A common sampling method is to discretize this ODE (e.g., via the explicit Euler method or a higher-order integrator) yielding a one-to-one mapping between successive time steps. For example, with a fixed step size $\Delta t$, the Euler scheme gives:
${\bm{x}}_{t-1}={\bm{x}}_{t}+\Delta t{\bm{v}}_{\theta}({\bm{x}}_{t},t,{\bm{c}})$ (8)
which produces a single, deterministic trajectory $\{{\bm{x}}_{T},{\bm{x}}_{T-1},\cdots,{\bm{x}}_{0}\}$ given any initial ${\bm{x}}_{T}$ and prompt ${\bm{c}}$.
This deterministic approach fails to meet GRPO’s policy update requirements in two key ways: (1) Eq. 6 requires computing $p({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$, but deterministic sampling cannot provide probabilities. This limitation affects not only GRPO but all policy gradient methods, including the basic online REINFORCE [55] policy gradient, which also require probability computation. (2) Exploration is crucial for RL performance. As shown in Section 5.3, reduced randomness significantly lowers training efficiency. Deterministic sampling, which lacks randomness beyond the initial seed, is especially problematic.
To address this limitation, we convert the deterministic Flow-ODE from Eq. 7 into an equivalent SDE that matches the original model’s marginal probability density function at all timesteps. We outline the key process here. A detailed proof is provided in Appendix A. Following [23, 28, 29], we construct a reverse-SDE formulation that preserves the marginal distribution:
@@
where $\mathrm{d}{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ control the level of stachasticity during generation. For rectified flow, the marginal score has a relation with velocity:
$\nabla\log p_{t}({\bm{x}})=-\frac{{\bm{x}}}{t}-\frac{1-t}{t}{\bm{v}}_{t}({\bm{x}}).$ (10)
Substituting Eq. 10 into Eq. 9 gives the final SDE:
@@
Eq. 12 reveals that the policy $\pi_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$ is an isotropic Gaussian distribution. We can easily compute the KL divergence between $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$ as a closed form:
$D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})=\frac{\|\overline{{\bm{x}}}_{t+\Delta t,\theta}-\overline{{\bm{x}}}_{t+\Delta t,\text{ref}}\|^{2}}{2\sigma_{t}^{2}\Delta t}=\frac{\Delta t}{2}\left(\frac{\sigma_{t}(1-t)}{2t}+\frac{1}{\sigma_{t}}\right)^{2}\|{\bm{v}}_{\theta}({\bm{x}}_{t},t)-{\bm{v}}_{\text{ref}}({\bm{x}}_{t},t)\|^{2}$ (13)
4.3 Denoising Reduction

### v3

@@ Abstract
We propose Flow-GRPO, the first method to integrate online reinforcement learning (RL) into flow matching models. Our approach uses two key strategies: (1) an ODE-to-SDE conversion that transforms a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) that matches the original model’s marginal distribution at all timesteps, enabling statistical sampling for RL exploration; and (2) a Denoising Reduction strategy that reduces training denoising steps while retaining the original number of inference steps, significantly improving sampling efficiency without sacrificing performance. Empirically, Flow-GRPO is effective across multiple text-to-image tasks. For compositional generation, RL-tuned SD3.5-M generates nearly perfect object counts, spatial relations, and fine-grained attributes, increasing GenEval accuracy from $63\%$ to $95\%$. In visual text rendering, accuracy improves from $59\%$ to $92\%$, greatly enhancing text generation. Flow-GRPO also achieves substantial gains in human preference alignment. Notably, very little reward hacking occurred, meaning rewards did not increase at the cost of appreciable image quality or diversity degradation.
@@ 1 Introduction
Flow matching [2, 3] models have become dominant in image generation [4, 5] due to their solid theoretical foundations and strong performance in producing high quality images. However, they often struggle with composing complex scenes involving multiple objects, attributes, and relationships [6, 7], as well as text rendering [8]. At the same time, online reinforcement learning (RL) [9] has proven highly effective in enhancing the reasoning capabilities of large language models (LLMs) [10, 11]. While previous research has mainly focused on applying RL to early diffusion-based generative models [12] and offline RL techniques like direct preference optimization [13] for flow-based generative models [14, 15], the potential of online RL in advancing flow matching generative models remains largely unexplored. In this study, we explore how online RL can be leveraged to effectively improve flow matching models.
@@
Training flow models with RL presents several critical challenges: (1) Flow models rely on a deterministic generative process based on ODEs [3], meaning they cannot sample stochastically during inference. In contrast, RL relies on stochastic sampling to explore the environment, learning by trying different actions and improving based on rewards. This need for stochasticity in RL conflicts with the deterministic nature of flow matching models. (2) Online RL depends on efficient sampling to collect training data, but flow models typically require many iterative steps to generate each sample, limiting efficiency. This issue is more pronounced with large models [5, 4]. To make RL practical for tasks like image or video generation, improving sampling efficiency is essential.
@@
To address these challenges, we propose Flow-GRPO, which integrates GRPO [16] into flow matching models for text-to-image (T2I) generation, using two key strategies. First, we adopt the ODE-to-SDE strategy to overcome the deterministic nature of the original flow model. By converting the ODE-based flow into an equivalent Stochastic Differential Equation (SDE) framework, we introduce randomness while preserving the original marginal distributions. Second, to improve sampling efficiency in online RL, we apply the Denoising Reduction strategy, which reduces denoising steps during training while keeping the full schedule during inference. Our experiments show that using fewer steps maintains performance while significantly reducing data generation costs.
@@
We evaluate Flow-GRPO on T2I tasks with various reward types. (1) Verifiable rewards, using the GenEval [17] benchmark and visual text rendering task. GenEval includes compositional image generation tasks (e.g., generating specific object counts, colors, and spatial relationships), which can be automatically assessed with object detection methods. Flow-GRPO improves the accuracy of Stable Diffusion 3.5 Medium (SD3.5-M) [4] from 63% to 95% on GenEval, outperforming the state-of-the-art GPT-4o [18] model. For visual text rendering, SD3.5-M’s accuracy increases from 59% to 92%, greatly enhancing its text generation ability. (2) Model-based rewards, such as the human preference Pickscore [19] reward. These results show that our framework is task independent, demonstrating its generalizability and robustness. Importantly, all improvements are achieved with very little reward hacking, as demonstrated in Figure 1.
@@
  • We are the first to introduce GRPO to flow matching models by converting deterministic ODE sampling into SDE sampling, showing the effectiveness of online RL for T2I tasks. Flow-GRPO improves SD3.5-M accuracy from 63% to 95% without noticeably compromising image quality.
  • We find that online RL for flow matching models does not require the standard long timesteps for training sample collection. By using fewer denoising steps during training and retaining the original steps during testing, we can significantly accelerate the training process.
  • We show that the Kullback-Leibler (KL) constraint effectively prevents reward hacking, where reward increases at the cost of image quality or diversity. KL regularization is not empirically equivalent to early stopping. With a proper KL term, we can match the high reward of the KL-free version while preserving image quality, albeit with longer training.
@@ 2 Related Work
Online RL has effectively improved the reasoning abilities of LLMs, such as DeepSeek-R1 [10] and OpenAI-o1 [11], using policy gradient methods like PPO [20] or value-free GRPO [16]. GRPO is more memory efficient by removing the need for a value network, so we adopt it in this work. PPO can also be applied to flow matching in a similar way.
@@
Diffusion models [21, 22, 23] add Gaussian noise to data and train a neural network to reverse the process. Sampling uses discrete DDPM steps or probability flow SDE solvers to generate high-fidelity outputs. Flow matching [2, 3] learns a continuous-time normalizing flow by directly matching the velocity field, allowing efficient deterministic sampling with only a few ODE steps. It achieves competitive FID with far fewer denoising steps than diffusion, making it the dominant choice in recent image [4, 5] and video [24, 25, 26, 27] generation models. Recent work [28, 29] unifies diffusion and flow models under an SDE/ODE framework. Our work builds on their theoretical foundations and introduces GRPO to flow-based models.
@@
Recent efforts to align pretrained T2I models with human preferences follow five main directions: (1) direct fine-tuning with scalar rewards [30, 31, 32, 33]; (2) Reward Weighted Regression (RWR) [34, 35, 36]; (3) Direct Preference Optimization (DPO) and variants [37, 38, 39, 14, 40, 41, 42, 43, 44, 36]; (4) PPO-style policy gradients [45, 46, 47, 48, 49, 50]; (5) training-free alignment methods [51, 52, 53]. These methods have successfully aligned T2I models with human preferences, improving aesthetics and semantic consistency. Building on this progress, we introduce GRPO for flow matching models, the backbone of today’s state-of-the-art T2I systems. Concurrent work [54] applies GRPO to text-to-speech flow models, but instead of converting the ODE to an SDE to inject stochasticity, they reformulate velocity prediction by estimating a Gaussian distribution (predicting both the mean and variance of velocity), which requires retraining the pre-trained model. Another study [55] also explores SDE-based stochasticity but focuses on inference-time scaling.
@@ 3 Preliminaries
In this section, we introduce the mathematical formulation of flow matching and describe how the denoising process can be mapped as a multi-step MDP.
Flow Matching.
Let ${\bm{x}}_{0}\sim X_{0}$ be a data sample from the true distribution, and ${\bm{x}}_{1}\sim X_{1}$ denote a noise sample. Recent advanced image-generation models (e.g., [4, 5]) and video-generation models (e.g., [24, 26, 25, 27]) adopt the Rectified Flow [3] framework, which defines the “noised” data ${\bm{x}}_{t}$ as
@@
Denoising as an MDP.
As shown in [12], the iterative denoising process in flow matching models can be formulated as a Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},\rho_{0},P,R)$. The state at step $t$ is ${\bm{s}}_{t}\triangleq({\bm{c}},t,{\bm{x}}_{t})$, the action is the denoised sample ${\bm{a}}_{t}\triangleq{\bm{x}}_{t-1}$ predicted by the model, and the policy is $\pi({\bm{a}}_{t}\mid{\bm{s}}_{t})\triangleq p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$. The transition is deterministic: $P({\bm{s}}_{t+1}\mid{\bm{s}}_{t},{\bm{a}}_{t})\triangleq(\delta_{\bm{c}},\delta_{t-1},\delta_{{\bm{x}}_{t-1}})$, and the initial state distribution is $\rho_{0}({\bm{s}}_{0})\triangleq(p({\bm{c}}),\delta_{T},\mathcal{N}(\mathbf{0},\mathbf{I}))$, where $\delta_{y}$ is the Dirac delta distribution centered at $y$. The reward is only given at the final step: $R({\bm{s}}_{t},{\bm{a}}_{t})\triangleq r({\bm{x}}_{0},{\bm{c}})$ if $t=0$, and 0 otherwise.
@@ 4 Flow-GRPO
In this section, we present Flow-GRPO, which enhances flow models using online RL. We begin by revisiting the core idea of GRPO [16] and adapting it to flow matching. We then show how to convert the deterministic ODE sampler into a SDE sampler with the same marginal distribution, introducing the stochasticity needed for applying GRPO. Finally, we introduce Denoise Reduction, a practical sampling strategy that significantly speeds up training without sacrificing performance.
GRPO on Flow Matching.
RL aims to learn a policy that maximizes the expected cumulative reward. This is often formulated as optimizing a policy $\pi_{\theta}$ with a regularized objective:
@@
Unlike other policy based methods like PPO [20], GRPO [16] provides a lightweight alternative, which introduces a group relative formulation to estimate the advantage.
Recall that the denoising process can be formulated as an MDP, as shown in Section 3. Given a prompt ${\bm{c}}$, the flow model $p_{\theta}$ samples a group of $G$ individual images $\{{\bm{x}}^{i}_{0}\}_{i=1}^{G}$ and the corresponding reverse-time trajectories $\{({\bm{x}}^{i}_{T},{\bm{x}}^{i}_{T-1},\cdots,{\bm{x}}^{i}_{0})\}_{i=1}^{G}$. Then, the advantage of the $i$-th image is calculated by normalizing the group-level rewards as follows:
@@
GRPO optimizes the policy model by maximizing the following objective:
$\mathcal{J}_{\text{Flow-GRPO}}(\theta)=\mathbb{E}_{{\bm{c}}\sim\mathcal{C},\{{\bm{x}}^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid{\bm{c}})}f(r,\hat{A},\theta,\varepsilon,\beta),$ (5)
@@
$f(r,\hat{A},\theta,\varepsilon,\beta)$ $=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg{(}\min\Big{(}r^{i}_{t}(\theta)\hat{A}^{i}_{t},\ \text{clip}\Big{(}r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\Big{)}\hat{A}^{i}_{t}\Big{)}-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Bigg{)},$
$r^{i}_{t}(\theta)$ $=\frac{p_{\theta}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}{p_{\theta_{\text{old}}}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}.$
From ODE to SDE.
GRPO relies on stochastic sampling in Eq. 4 and Eq. 5 to generate diverse trajectories for advantage estimation and exploration. Diffusion models naturally support this: the forward process adds Gaussian noise step by step, and the reverse process approximates a score-based SDE solver via a Markov chain with decreasing variance. In contrast, flow matching models use a deterministic ODE for the forward process:
@@
where ${\bm{v}}_{t}$ is learned via the flow matching objective in Eq. 2. A common sampling method is to discretize this ODE, yielding a one-to-one mapping between successive time steps.
This deterministic approach fails to meet the GRPO policy update requirements in two key ways: (1) $r^{i}_{t}(\theta)$ in Eq. 5 requires computing $p({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$, which becomes computationally expensive under deterministic dynamics due to divergence estimation. (2) More importantly, RL depends on exploration. As shown in Section 5.3, reduced randomness greatly lowers training efficiency. Deterministic sampling, with no randomness beyond the initial seed, is especially problematic.
To address this limitation, we convert the deterministic Flow-ODE from Eq. 6 into an equivalent SDE that matches the original model’s marginal probability density function at all timesteps. We outline the key process here. A detailed proof is provided in Appendix A. Following [23, 28, 29], we construct a reverse-time SDE formulation that preserves the marginal distribution:
@@
where $\mathrm{d}{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ control the level of stachasticity during generation. For rectified flow, Eq. 7 is specified as:
@@
Eq. 9 reveals that the policy $\pi_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$ is an isotropic Gaussian distribution. We can easily compute the KL divergence between $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$ in Eq. 5 as a closed form:
$D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})=\frac{\|\overline{{\bm{x}}}_{t+\Delta t,\theta}-\overline{{\bm{x}}}_{t+\Delta t,\text{ref}}\|^{2}}{2\sigma_{t}^{2}\Delta t}=\frac{\Delta t}{2}\left(\frac{\sigma_{t}(1-t)}{2t}+\frac{1}{\sigma_{t}}\right)^{2}\|{\bm{v}}_{\theta}({\bm{x}}_{t},t)-{\bm{v}}_{\text{ref}}({\bm{x}}_{t},t)\|^{2}$
Denoising Reduction.